Krepps Corporation produces a single product. Last year, Krepps manufactured 27,970 units and sold 22,200 units. Production cMultiple Choice o $221,949 o О $255,949 o $194,449 o О $254, 449 $264,449
Answer #1

Unit product cost under absorption costing

= (234,948+131,459+240,542+335,640)/27970

= 942,589/27970

= 33.70 per unit

Cost of ending inventory = (27970-22200)*33.70

= 194,449